OpenStudy (kittiwitti1):

Use the comparison property for bounding integrals to estimate the value of the integral... What is the comparison property, exactly?

OpenStudy (kittiwitti1):

\[\int_{1}^{5}(1+2x-\frac{1}{2}x^2)dx\]

OpenStudy (sshayer):

i do not know the comparison property but i can solve it.

OpenStudy (kittiwitti1):

Oh. But the question requires it I think? Let's see, maybe it is the same thing but labeled differently :p

OpenStudy (sshayer):

\[\int\limits_{1}^{5}\left( 1+2x-\frac{ x^2 }{ 2 } \right)dx=\left[ x+2\frac{ x^2 }{ 2 }-\frac{ x^3 }{ 2*3 } \right]~from ~1~\to~5\] \[=\left( 5-1 \right)+(5)^2-(1)^2-\frac{ 1 }{ 6 }\left\{ 5^3-1^3 \right\}=4+25-1-\frac{ 124 }{ 6 }=?\]

OpenStudy (kittiwitti1):

Oh, er... I don't think that's the property. But I will try nd see if it works. Sorry for the late response @sshayer thank you :)

OpenStudy (sshayer):

np